OpenStudy (anonymous):

Find an explicit rule for the nth term of the sequence. -5, -25, -125, -625, ...

OpenStudy (anonymous):

an = -5 • 5n an = 5 • -5n + 1 an = 5 • -5n an = -5 • 5n - 1

OpenStudy (aravindg):

this is a geometric progression with a=-5 r=5 cant u proceed from there?

OpenStudy (anonymous):

no, i have no idea how to do this type of problem :( @AravindG

OpenStudy (aravindg):

nth term of gp=ar^n-1 here -5(5)^(n-1)

OpenStudy (anonymous):

so the last option?

OpenStudy (anonymous):

@AravindG

OpenStudy (aravindg):

yep

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

what about Find an explicit rule for the nth term of the sequence. 9, 36, 144, 576, ... an = 9 • 4n - 1 an = 4 • 9n - 1 an = 9 • 4n an = 9 • 4n + 1